With the use of reduction of order for differential equations, reduce the following to first order and solve. (i) 𝑦 β€²β€² + 𝑒 𝑦𝑦 β€²3 = 0, (ii) π‘₯𝑦 β€²β€² + 2𝑦 β€² + π‘₯𝑦 = 0, 𝑦1 = sin π‘₯ π‘₯ , (iii) (1 βˆ’ π‘₯ 2 )𝑦 β€²β€² βˆ’ 2π‘₯𝑦 β€² + 2𝑦 = 0, 𝑦1 = π‘₯, (iv) 4π‘₯ 2𝑦 β€²β€² βˆ’ 3𝑦 = 0, 𝑦(1) = 3, 𝑦 β€² (1) = 2.5.Β 


1
Expert's answer
2021-11-29T14:17:17-0500

i)

"\ud835\udc66'' + \ud835\udc52^\ud835\udc66(\ud835\udc66')^3 = 0"

"y'=u"

"y''=uu'"


"uu'+e^yu^3=0"

"u'+e^yu^2=0"


"du\/u^2=-e^ydy"


"1\/u=e^y+c"


"\frac{}{}" "dx=(e^y+c)dy"


"x=e^y+cy+c_1"


ii)

"y''+p(x)y'+q(x)y=0"


"y''+2y'\/x+y=0"


"y_2=v(x)y_1=vsinx\/x"


"v(x)=c\int\frac{1}{y_1^2}e^{-\int p(x)dx}dx=c\int\frac{x^2}{sin^2x}e^{-2lnx}dx="


"=c\int\frac{1}{sin^2x}dx=c(-cotx)+k"


"v(x)=cotx"

"y_2=cotxsinx\/x=cosx\/x"


"y(x)=c_1sinx\/x+c_2cosx\/x"


iii)

"\ud835\udc66 ''\u2212 2\ud835\udc65\ud835\udc66 '\/(1 \u2212 \ud835\udc65 ^2 )+ 2\ud835\udc66\/(1 \u2212 \ud835\udc65^ 2 ) = 0"


"v(x)=c\int\frac{1}{y_1^2}e^{-\int p(x)dx}dx=c\int\frac{1}{x^2}e^{-ln(x^2-1)}dx=c\int\frac{1}{x^2(x^2-1)}dx="


"=c(\frac{ln\frac{x-1}{x+1}}{2}+\frac{1}{x})+k"


"v(x)=\frac{ln\frac{x-1}{x+1}}{2}+\frac{1}{x}"


"y_2=x(\frac{ln\frac{x-1}{x+1}}{2}+\frac{1}{x})=\frac{xln\frac{x-1}{x+1}}{2}+1"


"y(x)=c_1x+c_2(\frac{xln\frac{x-1}{x+1}}{2}+1)"


iv)

"y=x^k"

"(x^k)''=k(k-1)x^{k-2}"

"(4k^2-4k-3)x^k=0"

"4k^2-4k-3=0"

"k_1=-1\/2,k_2=3\/2"


"y_1=c_1/\sqrt x,y_2=c_2x^{3/2}"


"y(x)=y_1+y_2=c_1/\sqrt x+c_2x^{3/2}"


"y(1)=c_1+c_2=3"


"y'(x)=-c_1/(2x\sqrt x)+3c_2x^{1/2}/2"


"\ud835\udc66 '(1) =-c_1\/2+3c_2\/2= 2.5"


"c_2-3+3c_2=5"

"c_2=2,c_1=3-c_2=1"


"y(x)=1/\sqrt x+2x^{3/2}"